OpenStudy (anonymous):

Two legs of a triangle measures 7m and 13m. How many meters long is the hypotenuse? a.) 20 b.)6 c.)sqrt 218 d.)2 sqrt 30

OpenStudy (anonymous):

\[\sqrt{7^2+13^2}\]

OpenStudy (anonymous):

\[Pythagorean Theorem: a^2+b^2=c^2\] a = 7 or 13 (choose 1) b = 13 or 7 (choose the other 1) \[c=\sqrt{(7)^2+(13)^2} \]